# Quantum synchronization of twin limit-cycle oscillators

### v1

@@ Acknowledgements.
Here, we present some more detailed calculations and derivations, including the analysis of a single quantum trajectory of a twin limit cycle (TLC), the predictions of the locking behavior of TLCs based on their mean-field equations, the phase locking of standard quantum limit cycles, regions in TLC parameter space where the synchronization blockade exists, and the quantum mutual information of coupled TLCs.
I 1. Twin limit cycle and trajectory analysis
In this section, we examine the coexistence of limit cycles by analyzing the dynamics of a single trajectory in a quantum Liénard system. In Fig. S1, we specifically choose dissipation rates $\gamma_{i}$ that allow the effective potential $V(r)$ of the system to have two well-separated stable minima at $r_{1}=1$ (inner limit cycle) and $r_{2}=8$ (outer limit cycle), as well as an unstable maximum at $r_{c}=4$. The minima of $V(r)$ are sufficiently separated to facilitate the observation of the transition from one stable radius to the other due to intrinsic quantum noise. As shown in Fig. S1(a), the noise has to overcome a smaller potential difference, $\Delta V_{1}$, when transitioning from $r_{1}$ to $r_{2}$, in contrast to the larger potential difference, $\Delta V_{2}$, for the reverse direction. This asymmetry in the potential differences is evident in the time evolution of a single trajectory presented in Fig. S1(b), where the system spends more time in the outer limit cycle at $r_{2}$ compared to the inner limit cycle at $r_{1}$. Notably, this behavior of the quantum trajectories is independent of the choice of the initial state. The steady-state density matrix can be understood as the long-time average of many quantum trajectories, which results in the two ring-like structures found in the corresponding Wigner function, as shown in Fig. S1(c).
@@
II 2. Approximation of the mean-field equations
@@
II.1 2.1. Limit cycles of the same type: (in,in) and (out,out)
@@
II.2 2.2. Limit cycles of different type: (in,out) and (out,in)
@@
III 3. Phase locking of single standard limit-cycle oscillators
III.1 3.1. Single Driven Standard Limit Cycle
@@
Figure S2: Phase locking of standard limit cycles. Solid curves denote the maxima of $P_{1}$ whereas dashed curves represent $\text{arg}(\langle\tilde{a}_{j}\rangle)$. (a), (c) $K_{j}=0$. (b), (d) $\Delta_{j}=0$. In panels (a) and (b), we set the drive strength to $\Omega_{j}=\gamma^{A}_{1}/6$ and the gain and damping rates to $\gamma^{A}_{1}=\gamma^{B}_{1}$, $\gamma^{A}_{2}=2.5\gamma^{A}_{1}$, and $\gamma^{B}_{2}=0.1\gamma^{A}_{1}$. In panels (c) and (d), we set the drive strength to $\Omega_{j}=\gamma^{A}_{1}/6$ and the gain and damping rates to $\gamma^{A}_{3}=\gamma^{B}_{3}$, $\gamma^{A}_{4}=3.75\gamma^{A}_{3}$, and $\gamma^{B}_{4}=0.15\gamma^{A}_{3}$.
@@
III.2 3.2. Two Coupled Standard Limit Cycles
@@
For $\delta=K=0$, we expect that the relative phase locks to $\phi_{BA}\in\{0,\pi\}$ for $\gamma^{A}_{2}=\gamma^{B}_{2}$ and to $\phi_{BA}=\pi/2$ ($\phi_{BA}=-\pi/2$) for $\gamma^{A}_{2}>\gamma^{B}_{2}$ ($\gamma^{A}_{2}<\gamma^{B}_{2}$). These three cases coincide with the mean-field predictions for TLCs Eqs. S6 and S11, where $(\mathrm{in,in})$ and $(\mathrm{out,out})$ correspond to $\gamma^{A}_{2}=\gamma^{B}_{2}$, $(\mathrm{in,out})$ to $\gamma^{A}_{2}>\gamma^{B}_{2}$, and $(\mathrm{out,in})$ to $\gamma^{A}_{2}<\gamma^{B}_{2}$.
IV 4. Quantum Synchronization Blockade
@@
In a system of two coupled TLC oscillators, we vary the rates $\gamma^{A}_{2}$, $\gamma^{A}_{3}$, and $\gamma^{A}_{4}$ individually by keeping $\gamma^{A}_{1}=\gamma^{B}_{1}$, $\gamma^{B}_{2}=2.5\gamma^{A}_{1}$, $\gamma^{B}_{3}=1.04\gamma^{A}_{1}$, and $\gamma^{B}_{4}=0.096\gamma^{A}_{1}$ fixed. The resulting blockades are illustrated in Figs. S3(b) to S3(g) and exist in a narrower range of $\gamma^{A}_{j}/\gamma^{B}_{j}\in[0.97,1.03]$ compared to the system of two standard limit cycles presented in Fig. S3(a). Thus, the blockade in a pair of TLCs is more susceptible to variations in gain and damping rates than that of two standard limit cycles.
Figure S3: Regions of the quantum synchronization blockade for $g=\gamma^{A}_{1}$, $\gamma^{A}_{1}=\gamma^{B}_{1}$, and $\gamma^{B}_{2}=2.5\gamma^{A}_{1}$. In all panels, maxima in $\phi_{BA}$ are denoted by dotted lines. (a) Two standard limit cycles with $\gamma^{j}_{3}=\gamma^{j}_{4}=0$ and $\gamma^{B}_{2}=2.5\gamma^{A}_{1}$. The blockade region is limited by the two vertical lines and indicated by the arrow (omitted in panels (b)–(g)). The color scale is linear in the interval $[-10^{-3},10^{-3}]$ and logarithmic elsewhere. (b)–(d) $P_{2}^{\mathrm{in},\mathrm{in}}$. (e)–(g) $P_{2}^{\mathrm{out},\mathrm{out}}$. In panels (b)–(g), the higher-order rates are set to $\gamma^{B}_{3}=1.04\gamma^{A}_{1}$ and $\gamma^{B}_{4}=0.096\gamma^{A}_{1}$ as in the main text. The color scale is linear in the interval $[-10^{-6},10^{-6}]$ and logarithmic elsewhere.
V 5. Quantum Mutual Information

### v2

@@ Acknowledgements.
Here, we present some more detailed calculations and derivations, including the analysis of a single quantum trajectory of a twin limit cycle (TLC), the frequency synchronization of a TLC, the predictions of the locking behavior of TLCs based on their mean-field equations, the phase locking of standard quantum limit cycles, regions in TLC parameter space where the synchronization blockade exists, and the quantum mutual information of coupled TLCs.
I 1. Twin limit cycle and quantum trajectory analysis
In this section, we examine the coexistence of limit cycles by analyzing the dynamics of a single quantum trajectory in a quantum Liénard system. In Fig. S1, we specifically choose dissipation rates $\gamma_{i}$ that allow the effective potential $V(r)$ of the system to have two well-separated stable minima at $r_{1}=1$ (inner limit cycle) and $r_{2}=8$ (outer limit cycle), as well as an unstable maximum at $r_{c}=4$. The minima of $V(r)$ are sufficiently separated to facilitate the observation of the transition from one stable radius to the other due to intrinsic quantum noise. As shown in Fig. S1(a), the noise has to overcome a smaller potential difference, $\Delta V_{1}$, when transitioning from $r_{1}$ to $r_{2}$, in contrast to the larger potential difference, $\Delta V_{2}$, for the reverse direction. This asymmetry in the potential differences is evident in the time evolution of a single trajectory presented in Fig. S1(b), where the system spends more time in the outer limit cycle at $r_{2}$ compared to the inner limit cycle at $r_{1}$. Notably, this behavior of the quantum trajectories is independent of the choice of the initial state. The steady-state density matrix can be understood as the long-time average of many quantum trajectories, which results in the two ring-like structures found in the corresponding Wigner function, as shown in Fig. S1(c).
@@
II 2. Frequency Synchronization
In this section, we discuss the frequency synchronization of TLCs and compare it with the phase synchronization measure. We utilize the power spectrum to analyze the frequency entrainment of the TLC, which is defined as follows
$S_{\alpha}(\omega)=\lim_{t\to\infty}\int\limits_{-\infty}^{\infty}\mathrm{d}\tau\langle a^{\dagger}_{\alpha}(t+\tau)a_{\alpha}(t)\rangle\mathrm{e}^{-\mathrm{i}\omega\tau}\,$, (S1)
where $\alpha=\text{in},\text{out}$. The power spectrum defined above is based on two-time correlations of the truncated annihilation operators, which we define as
[…]
Note the difference to Eq. 7, where the factor $\sqrt{n+1}$ is not included.
Figure S2: Phase and frequency synchronization of a driven TLC with $\gamma_{2}=2.5\gamma_{1}$, $\gamma_{3}=1.04\gamma_{1}$, and $\gamma_{4}=0.096\gamma_{1}$. (a), (b) Arnold tongues for inner and outer limit cycles as a function of detuning $\Delta$ and drive strength $\Omega$ describing the magnitude of phase synchronization. The black curves denote contour lines at $5$. (c) Power spectra $S_{\alpha}(\omega)$ corresponding to the red dot at $\Delta=0$ and $\Omega=0.25\gamma_{1}$ in panels (a) and (b). (d) Power spectra $S_{\alpha}(\omega)$ corresponding to the red triangle at $\Delta=0$ and $\Omega=4.5\gamma_{1}$ in panels (a) and (b). Both spectra exhibit a peak at $\omega=0$ due to injection locking. The power spectra are based on the truncated annihilation operators defined in Eq. S2.
We start by discussing the phase synchronization measure, i.e., the maximum of $P_{1}^{\alpha}(\phi)$ defined in Eq. 8. The phase synchronization measure shows Arnold tongues for both the inner and outer limit cycle as functions of detuning $\Delta$ and drive strength $\Omega$, as illustrated in Figs. S2(a) and S2(b). Both limit cycles exhibit a vanishing amount of synchronization for lower $\Omega$ values and require higher coupling strength to observe a large amount of synchronization due to the presence of higher-order dissipators.
The power spectrum is well-defined for both the limit cycles for small values of $\Omega$. One such example is shown in Fig. S2(c) where $\Omega=0.25\gamma_{1}$ and $\Delta=0$. However, no frequency locking occurs when $\Delta\neq 0$ for such smaller drive strength. In regions of stronger drive, where a significant degree of phase synchronization is observed, the spectrum of the inner limit cycles broadens considerably, as illustrated in Fig. S2(d) for $\Omega=4.5\gamma_{1}$ and $\Delta=0$. Such spectral broadening results from changes in the population distribution within the inner limit cycle, shifting toward the critical radius for higher $\Omega$ values. With a further increase in $\Omega$, the population increasingly overlaps with the outer limit cycle, expected to lead to further broadening of the power spectrum. This shift hampers the quantization of synchronization within the inner limit cycle when analyzed through the power spectrum, limiting its applicability to a narrow region that holds little significance. Therefore, the phase synchronization measure is more sensitive and applicable even at higher drive strengths for the TLC oscillators.
III 3. Expansion of the mean-field equations
@@
III.1 3.1. Limit cycles of the same type: (in,in) and (out,out)
@@
III.2 3.2. Limit cycles of different type: (in,out) and (out,in)
@@
IV 4. Phase locking of single standard limit-cycle oscillators
IV.1 4.1. Single Driven Standard Limit Cycle
@@
Figure S3: Phase locking of standard limit cycles. Solid curves denote the maxima of $P_{1}$ whereas dashed curves represent $\text{arg}(\langle\tilde{a}_{j}\rangle)$. (a), (c) $K_{j}=0$. (b), (d) $\Delta_{j}=0$. In panels (a) and (b), we set the drive strength to $\Omega_{j}=\gamma^{A}_{1}/6$ and the gain and damping rates to $\gamma^{A}_{1}=\gamma^{B}_{1}$, $\gamma^{A}_{2}=2.5\gamma^{A}_{1}$, $\gamma^{B}_{2}=0.1\gamma^{A}_{1}$, and $\gamma^{j}_{3}=\gamma^{j}_{4}=0$. In panels (c) and (d), we set the drive strength to $\Omega_{j}=\gamma^{A}_{3}/6$ and the gain and damping rates to $\gamma^{j}_{1}=\gamma^{j}_{2}=0$, $\gamma^{A}_{3}=\gamma^{B}_{3}$, $\gamma^{A}_{4}=3.75\gamma^{A}_{3}$, and $\gamma^{B}_{4}=0.15\gamma^{A}_{3}$.
@@
IV.2 4.2. Two Coupled Standard Limit Cycles
@@
For $\delta=K=0$, we expect that the relative phase locks to $\phi_{BA}\in\{0,\pi\}$ for $\gamma^{A}_{2}=\gamma^{B}_{2}$ and to $\phi_{BA}=\pi/2$ ($\phi_{BA}=-\pi/2$) for $\gamma^{A}_{2}>\gamma^{B}_{2}$ ($\gamma^{A}_{2}<\gamma^{B}_{2}$). These three cases coincide with the mean-field predictions for TLCs Eqs. S8 and S13, where $(\mathrm{in,in})$ and $(\mathrm{out,out})$ correspond to $\gamma^{A}_{2}=\gamma^{B}_{2}$, $(\mathrm{in,out})$ to $\gamma^{A}_{2}>\gamma^{B}_{2}$, and $(\mathrm{out,in})$ to $\gamma^{A}_{2}<\gamma^{B}_{2}$.
V 5. Quantum Synchronization Blockade
@@
In a system of two coupled TLC oscillators, we vary the rates $\gamma^{A}_{2}$, $\gamma^{A}_{3}$, and $\gamma^{A}_{4}$ individually by keeping $\gamma^{A}_{1}=\gamma^{B}_{1}$, $\gamma^{B}_{2}=2.5\gamma^{A}_{1}$, $\gamma^{B}_{3}=1.04\gamma^{A}_{1}$, and $\gamma^{B}_{4}=0.096\gamma^{A}_{1}$ fixed. The resulting blockades are illustrated in Figs. S4(b) to S4(g) and exist in a narrower range of $\gamma^{A}_{j}/\gamma^{B}_{j}\in[0.97,1.03]$ compared to the system of two standard limit cycles presented in Fig. S4(a). Thus, the blockade in a pair of TLCs is more susceptible to variations in gain and damping rates than that of two standard limit cycles.
Figure S4: Regions of the quantum synchronization blockade for $g=\gamma^{A}_{1}$, $\gamma^{A}_{1}=\gamma^{B}_{1}$, and $\gamma^{B}_{2}=2.5\gamma^{A}_{1}$. In all panels, maxima in $\phi_{BA}$ are denoted by dotted lines. (a) Two standard limit cycles with $\gamma^{j}_{3}=\gamma^{j}_{4}=0$ and $\gamma^{B}_{2}=2.5\gamma^{A}_{1}$. The blockade region is limited by the two vertical lines and indicated by the arrow (omitted in panels (b)–(g)). The color scale is linear in the interval $10^{-3},10^{-3}]$ and logarithmic elsewhere. (b)–(d) $P_{2}^{\mathrm{in},\mathrm{in}}$. (e)–(g) $P_{2}^{\mathrm{out},\mathrm{out}}$. In panels (b)–(g), the higher-order rates are set to $\gamma^{B}_{3}=1.04\gamma^{A}_{1}$ and $\gamma^{B}_{4}=0.096\gamma^{A}_{1}$ as in the main text. The color scale is linear in the interval $10^{-6},10^{-6}]$ and logarithmic elsewhere.
VI 6. Quantum Mutual Information